OpenStudy (anonymous):

calculate the average speed for a round trip with a “going” average speed of 30 miles per hour and a “returning” average speed of 60 miles per hour

OpenStudy (zarkon):

40

OpenStudy (anonymous):

could you please explain :) @zarkon

OpenStudy (anonymous):

... for potential readers

OpenStudy (anonymous):

@Australopithecus : your answer is wrong

OpenStudy (anonymous):

Let 2d be the round trip distance. Average speed would be 2d divided by the time to travel a distance of 2d.\[t=\frac{d}{r} \]The time out is\[t=\frac{d}{30} \]and\[t=\frac{d}{60} \] is the time to return.\[\frac{2d}{\frac{d}{30}+\frac{d}{60}}=40 \text{ mph} \]

OpenStudy (anonymous):

@robtobey :)

OpenStudy (anonymous):

this is one of the tricky problems ....